OpenStudy (anonymous):

Solve 6n + 2 - 3n = 17.

OpenStudy (anonymous):

6n+2-3n=17 (6-3)n+2=17 3n+2=17 3n+2-2=17-2 3n=15 3n/3=15/3 n=5